OpenStudy (anonymous):

Isaac is cutting wood to build a clock. He wants the clock to have a long hand that extends from the center to the edge of the clock face. What must the approximate area of the face be if the hand he is using is 8 inches long? Use three point one four for pi.

OpenStudy (anonymous):

Okay. This geometry question is pretty simple. It's all about finding out the area of a circle. I'll walk you through it, kay? 1. We know that clocks are circles. And the question is asking us, "What must be the approximate area of the face (of the clock) be?" That's why we're gonna find the area of a circle. 2. We have to know the formula for finding the area of a circle. Area of a circle = π * r^2 (That's pi, times the radius, after the radius has been squared.) 3. Okay, so let's do it! First, we know the radius is 8. (You know what a radius is, right? The radius is the length from the center of a circle, all the way out into the edge. I'll attach a picture of a radius of a circle.) 4. We know that Pi = 3.14 ***Take one more look at the formula for finding the are of a circle.*** Area = π * r^2 5. What's the radius, squared? 8 * 8 = 64. 6. Now, what's 64 * 3.14? 64 * 3.14 = 200.96 That's your answer! The area = 200.96 inches squared!